OpenStudy (anonymous):

Use the Distributive Property and simplify the expression. 6 + 4(3x – 2) + 5

OpenStudy (johnweldon1993):

Distributive.....so take that 4 and distribute that into the parenthesis (multiply it by each term in the parenthesis) what do you have after that?

OpenStudy (amigatour):

Remember the rules of PEMDAS. However, you got nothing to solve in the parenthesis, so move on on multiply 4(3x-2) and then simplify.

OpenStudy (shamil98):

Please refrain from giving the answer, @charlotte123 .

OpenStudy (charlotte123):

You need the Distributive Property and PEMDAS for this one ~

OpenStudy (charlotte123):

Now now Sham =.=

OpenStudy (therealmeeeee):

6 + 4(3x-2) + 5 First, multiply 4 to (3x-2) 6 + 4(3x-2) + 5 = 6 + 12x - 8 + 5 Next, group similar terms and compute them = [12x] + [6 - 8 +5] = 12x + 3 (D)

OpenStudy (therealmeeeee):

rhe answer is D

OpenStudy (charlotte123):

There are no choices xD

OpenStudy (charlotte123):

A test? Oh oh.............I'm just gonna step away...........real slowly.........*runs* O_O

OpenStudy (therealmeeeee):

X_X lol @charlotte123

OpenStudy (charlotte123):

I smell a mod coming on......RUN FOR YOUR LIFE MAN

OpenStudy (therealmeeeee):

i already am

OpenStudy (anonymous):

WHATS GOING ON